What is the magnification of a telescope which has an objective lens of focal length 2.50 m and a diameter of 1.00 m, and an eyepiece of focal length 4.50 cm?

What is the angular resolution of this telescope for visible light (wavelength = 550 nm)?

What is the length of the telescope?
Answer #1

fo = objective lens focal length = 2.50 m

fe = eyepiece focal length = 4.50 cm = 0.045 m

magnification is given as

m = - fo/fe = - 2.5 / 0.045 = - 55.6

angular resolution = 1.22 /d

angular resolution = 1.22 (550 x 10-9) / (1) = 6.71 x 10-7 rad

Length of telescope =L = fo + fe = 2.5 + 0.045 = 2.545 m